a stone is projected from the ground with a velocity of 14ms-1.one second later it clears a wall 2m high .the angle of projection is (g= 10ms-2).

Grade:11

2 Answers

Zameer Ansari
6 years ago
H=2m g= -10m/s^2 velocity in upward direction is unknown=ut=1sH=ut+1/2at^22=u×1+1/2×(-10)×1u=7m/sNow, as we know velocity upwards, u=v.sinx where v is the given velocity in question..7=14.sinxSinx=1/2X=30°
Zameer Ansari
6 years ago
H=2m g= -10m/s^2 velocity in upward direction is unknown=u t=1s H=ut+1/2at^2 2=u×1+1/2×(-10)×1 u=7m/s Now, as we know velocity upwards,u u=v.sinx where v is the given velocity in question.. 7=14.sinx Sinx=1/2 X=30°